### v1

# An Efficient Scene Coordinate Encoding and Relocalization Method

Abstract
Scene Coordinate Regression (SCR) is a visual localization technique that utilizes deep neural networks (DNN) to directly regress 2D-3D correspondences for camera pose estimation. However, current SCR methods often face challenges in handling repetitive textures and meaningless areas due to their reliance on implicit triangulation. In this paper, we propose an efficient scene coordinate encoding and relocalization method. Compared with the existing SCR methods, we design a unified architecture for both scene encoding and salient keypoint detection, enabling our system to focus on encoding informative regions, thereby significantly enhancing efficiency. Additionally, we introduce a mechanism that leverages sequential information during both map encoding and relocalization, which strengthens implicit triangulation, particularly in repetitive texture environments. Comprehensive experiments conducted across indoor and outdoor datasets demonstrate that the proposed system outperforms other state-of-the-art (SOTA) SCR methods. Our single-frame relocalization mode improves the recall rate of our baseline by 6.4% and increases the running speed from 56 to 90. Furthermore, our sequence-based mode increases the recall rate by 11% while maintaining the original efficiency.
I Introduction
Visual relocalization is crucial for many applications such as mobile robots and augmented reality, offering a practical balance between cost and accuracy [1]. Its core task involves estimating the camera pose using an existing map, which can be broadly divided into two categories: explicit maps and implicit maps. Explicit maps consist of a 3D point cloud combined with visual descriptors of 2D features [2, 3], while implicit maps encode the scene using a neural network [1, 4]. Feature matching (FM)-based methods construct explicit maps using techniques like structure-from-motion (SFM) [3, 5] or simultaneous localization and mapping (SLAM) [2, 6]. During relocalization, these methods rely on image retrieval and feature matching to establish 2D-3D correspondences between pixels in the query image and 3D points in the map. However, this approach often suffers from large map sizes due to the storage of visual features and 3D models. Additionally, the process can be computationally inefficient, as the feature matching needs to be performed between the query image and multiple candidates.
In contrast, scene coordinate regression (SCR) methods encode map information into a neural network [1, 7]. They directly regress the 3D coordinates in the world coordinate system correspondences to the image pixels, eliminating the need for explicit 2D-3D corresponding searching. Compared to FM-based methods, SCR methods enable faster relocalization with a smaller map. Recently, ACE [8] proposes a generalized convolutional neural network (CNN) to extract feature maps for scene encoding. Building on this pre-trained backbone, it utilizes a compact 4MB-sized multilayer perceptron (MLP) head as the map, enabling new scene encoding in five minutes only without requiring a 3D model or depth information. This approach achieves state-of-the-art accuracy and efficiency.
Figure 1: In this paper, we propose an efficient SCR system that leverages deep neural networks for both mapping and relocalization. Our system delivers accurate and efficient relocalization, while maintaining fast mapping and a compact map size, achieving state-of-the-art performance.
Despite their advantages, SCR methods face several notable limitations. First, the presence of many texture-less and non-informative regions in the image significantly increases the map encoding errors [9]. Second, existing RGB-only SCR methods rely heavily on implicit triangulation, i.e., the neural network must consistently identify and map the same regions across different images to accurate 3D points. This reliance causes a sharp decline in performance in environments with repetitive textures [10]. Third, during relocalization, all pixel-to-3D matches are considered for pose estimation since the reliability of specific regions is unknown, which significantly reduces efficiency. Efforts have been made to address these challenges. FocusTune [9] mitigates encoding errors by focusing on informative regions but requires a 3D model of the scene, which limits its practicality. For the second issue, GLACE [10] introduces global image descriptors to better distinguish similar regions across images, yet this approach comes at the cost of reduced mapping and localization efficiency.
In this paper, we present a novel SCR system designed to tackle these challenges. To address the first and third issues, we introduce a unified and scene-agnostic CNN backbone that simultaneously encodes the scene and identifies salient regions, focusing mapping and relocalization efforts exclusively on these informative areas. For the second challenge, we explicitly associate the salient regions across different images, leveraging sequential information to enhance implicit triangulation and relocalization. In summary, the main contributions of this work are as follows:
  • We introduce a unified, scene-agnostic CNN backbone for scene encoding and salient region detection, which mitigates accuracy degradation caused by non-informative areas.
  • We propose a novel method that leverages sequential information to strengthen the implicit triangulation in SCR systems, improving mapping accuracy, particularly in environments with repetitive textures.
  • We design two relocalization modes: a single-frame mode that delivers one-shot relocalization with high efficiency and a sequence mode that utilizes temporal information to enhance relocalization performance.
  • Extensive experimental results show that our methods achieve substantial improvement in both efficiency and effectiveness. We provide the source code at https://github.com/sair-lab/SeqACE.
II Related Works
II-A Feature Matching Based Relocalization
FM-based visual relocalization methods remain the state-of-the-art for large-scale and challenging environments [10]. These approaches rely on 3D models generated through SfM or SLAM [2, 3, 11, 12]. During relocalization, candidate map images similar to the query image are retrieved using techniques such as bag-of-words (BoW) or image retrieval methods [3, 6]. Feature matching is then performed between the query image and each retrieved map image. Since the keypoints in the map images have already been triangulated, 2D-3D correspondences between the pixels in the query image and the 3D points in the point cloud model can be established. Finally, the 6-DOF pose of the query image is estimated using geometric techniques, such as Perspective-n-Point (PnP) and bundle adjustment [13, 14].
However, FM-based methods require storing 2D keypoints, visual descriptors, and a 3D point cloud, resulting in relatively large map sizes. For large-scale scenes, managing these maps demands significant memory resources. Furthermore, during relocalization, multiple candidate frames are typically retrieved for feature matching, which imposes computational overhead and reduces efficiency.
II-B Absolute Pose Regression
Absolute pose regression (APR) approaches use a DNN to regress camera poses from RGB images directly. They have also garnered notable attention recently due to their good efficiency and simple architecture. PoseNet [15] is the first ARP solution which uses GoogLeNet [16] as the backbone. The following works explore different architectural designs to enhance ARP systems [17, 18]. Other works leverage sequential information to improve APR accuracy by incorporating temporal filtering [19] or visual odometry [20]. Recent ARP works utilize novel synthesis techniques to generate high-quality training data [21] or integrate them for fine-tuning [22]. However, ARP approaches still fall short in accuracy compared to SCR and FM-based methods.
II-C Scene Coordinate Regression
Different from ARP approaches, SCR methods regress the corresponding 3D coordinates in the scene for 2D pixels in the query image. Most existing SCR methods require depth information or a 3D model of the scene. Dong et al. [7] utilize a pre-learned feature extractor and scene region classifier to encode a scene with few-shot images, achieving efficient mapping. However, the additional modules, e.g., feature detection and hierarchical classification, limit its localization speed to only $5$ even on a powerful GPU. Zhou et al. [23] propose KFNet. They leverage the sequential information by integrating the traditional Kalman filter [24] into the SCR pipeline. Bui et al. [25] propose D2S, which builds on a pre-trained sparse feature extraction network. They establish a mapping between feature descriptors and 3D scene points.
DSAC++ [26] is the first SCR method that does not rely on depth or 3D model information. However, it requires a long training time. DSAC* [1] simplifies the DSAC++ training procedure by unifying two separate initialization steps, reducing the training time from 6 days to 15 hours. ACE [8] replaces the backbone of DSAC* with a pre-trained CNN, requiring only a new header to be trained when encoding a new scene, which greatly improves the mapping efficiency. This reduces the mapping time to 4 minutes and the map size to 4MB. FocusTune [9] and GLACE [10] are variations of ACE. FocusTune projects a 3D model onto training images to avoid encoding meaningless image regions, but this introduces the dependency on the 3D model. GLACE incorporates global image descriptors into the ACE pipeline to enhance the distinguishability of similar regions.
III Preliminaries
III-A Problem Statement
The objective is to estimate the pose $\mathbf{T}_{wc}$ of a given image $I$, where $\mathbf{T}_{wc}$ denotes the transformation from the camera coordinate system to the world coordinate system. Given a set of 2D-3D correspondences $\mathcal{C}={(\mathbf{p}_{i},\mathbf{P}_{i}^{w})}$, where $\mathbf{p}_{i}$ represents a pixel in the image $I$ and $\mathbf{P}_{i}^{w}$ is its corresponding 3D point in the world coordinate system, the pose can be efficiently computed using PnP algorithms or iterative optimization techniques. To obtain these correspondences $\mathcal{C}$, SCR methods aim to learn a mapping from 2D pixels to their corresponding 3D points:
$\mathbf{P}_{i}^{w}=f\left(\mathcal{P}\left(\mathbf{p}_{i},I\right);\mathbf{w}\right),$ (1)
where $\mathcal{P}\left(\mathbf{p}_{i},I\right)\in\mathbb{R}^{{C_{p}}\times{H_{p}}\times{W_{p}}}$ is a small patch centered at $\mathbf{p}_{i}$ in the image $I$, and $f$ is a neural network with learnable parameters $\mathbf{w}$. Existing SCR pipelines usually take grayscale inputs with $C_{p}=1$. As there is no explicit patch extraction, the patch size is determined by the receptive field of the neural network. Considering the trade-off between invariance and discriminative power, the patch size is typically set to $H_{p}=W_{p}=81\text{px}$ in many existing architectures [8, 10].
III-B Accelerated Coordinate Encoding
Training $f$ in Eq. 1 using only posed images takes a long time [1]. To accelerate this process, Accelerated Coordinate Encoding (ACE) [8] proposes to split $f$ into a scene-agnostic backbone $f_{B}$ and a scene-specific MLP head $f_{H}$:
$\mathbf{P}_{i}^{w}=f_{H}\left(\mathbf{f}_{i};\mathbf{w}_{H}\right),\;\text{with}\ \mathbf{f}_{i}=f_{B}\left(\mathcal{P}\left(\mathbf{p}_{i},I\right);\mathbf{w}_{B}\right),$ (2)
where $f_{B}$ predicts a high-dimensional feature vector $\mathbf{f}_{i}$ for each patch, and $f_{H}$ maps $\mathbf{f}_{i}$ to the corresponding 3D point:
$f_{B}:\mathbb{R}^{{C_{p}}\times{H_{p}}\times{W_{p}}}\xrightarrow{}\mathbb{R}^{C_{\mathbf{f}}},\;\text{and}\ f_{H}:\mathbb{R}^{C_{\mathbf{f}}}\xrightarrow{}\mathbb{R}^{3},$ (3)
where $C_{\mathbf{f}}$ is the dimension of $\mathbf{f}_{i}$. The backbone $f_{B}$ is pre-trained on a large dataset. When mapping a new scene, only the MLP head $f_{H}$ needs to be trained while $f_{B}$ remains fixed. Thus, the MLP head serves as the map. This approach significantly reduces both the map size and the mapping time.
Figure 2: Our system comprises a scene-agnostic module and a scene-specific module. The scene-agnostic module identifies informative image patches and generates high-dimensional features for each patch. The scene-specific module, implemented using 1D convolutional layers, predicts the corresponding 3D scene point for each patch. These 2D-3D correspondences are then utilized for relocalization.
IV Methods
IV-A Overview
The overall structure of the proposed system is depicted in Fig. 2. It comprises two main components: a scene-agnostic module and a scene-specific module. The scene-agnostic module generates feature vectors and saliency scores for image patches, while the scene-specific module, designed as an MLP head with several 1D convolutional layers, maps these feature vectors to 3D scene points. The detailed design of the scene-agnostic module is presented in Section IV-B. Section IV-C describes how sequential information is utilized to enhance the training of the scene-specific module, and Section IV-D outlines our relocalization methods.
IV-B Scene-Agnostic Module
In this section, we aim to design an efficient and scene-agnostic module capable of (1) detecting informative image patches and (2) mapping these patches into high-dimensional features for scene encoding. As many learning-based patch and keypoint detection methods have been proposed [3, 27], a straightforward solution is to utilize them directly for salient region detection along with the ACE backbone. However, employing two separate deep neural networks would significantly reduce online efficiency. Inspired by [3, 6], we think that these two tasks can be accomplished using a shared backbone. To achieve this, we extend the ACE backbone by adding a lightweight keypoint detection head (KDH). The KDH consists of just three convolutional layers, introducing trivial computational overhead. Its input is the feature map from the backbone, and its output is a tensor $\mathcal{T}_{K}\in\mathbb{R}^{65\times\frac{H}{8}\times\frac{W}{8}}$, where $H$ and $W$ denote the height and width of the input image, respectively. The 65 channels correspond to an $8\times 8$ grid region and a dustbin channel representing the absence of keypoints. This tensor is processed through a softmax layer and resized into a $H\times W$ heatmap, which scores the saliency of the corresponding image pixels. Since the KDH, like the backbone $f_{B}$ in Eq. 2, is also scene-agnostic, we combine them into a unified module denoted as $f_{S}$, with its parameters represented by $\mathbf{w}_{S}$. Then the second equation in Eq. 2 is updated to:
$\mathbf{f}_{i},\mathcal{S}_{i}=f_{S}\left(\mathcal{P}\left(\mathbf{p}_{i},I\right);\mathbf{w}_{S}\right),$ (4)
where $\mathcal{S}_{i}$ is the salient score of $\mathbf{p}_{i}$.
We train $f_{S}$ using a two-step pipeline on the ScanNet dataset [28]. In the first step, we follow the approach outlined in [8] to train the backbone $f_{B}$. In the second step, the parameters of $f_{B}$ are fixed, and we utilize distillation techniques to train the KDH. Theoretically, any keypoint detector could serve as the teacher network. For our setup, we use the prediction of SuperPoint [27] as supervision signals. The loss function is defined as:
$\mathcal{L}_{K}=\mathcal{L}_{CE}\left(\mathcal{T}_{K},\mathcal{T}^{\prime}_{K}\right),$ (5)
where $\mathcal{L}_{CE}$ is the cross-entropy loss and $\mathcal{T}^{\prime}_{K}$ is the prediction of the teacher network. We also experimented with a multi-task learning approach [3] to train $f_{S}$, but found that the two-step pipeline produced better results. This is probably because the difficulty of the two tasks differs significantly: specifically, training a keypoint detector is much simpler than training a universal scene encoder. This disparity makes it challenging to balance the two tasks effectively.
Figure 3: Sampling strategies comparison. ACE randomly samples image patches. FocusTune performs uniform sampling around the projected points of the 3D model. We use keypoint detection to select repeatable image patches.
IV-C Sequence-Based Mapping
Saliency-Based Buffer Sampling
Buffer sampling is the key to fast mapping. Since the backbone, i.e., $f_{B}$ or $f_{S}$, is scene-agnostic, it only needs to be executed once when mapping a new scene. The resulting outputs, $\{\mathbf{f}_{i}\}$, can be cached and subsequently used to train $f_{H}$. To enhance efficiency, not all image patches are utilized during training. In ACE [8], 1024 image patches are randomly sampled per image, and their corresponding $\mathbf{f}_{i}$, $\mathbf{p}_{i}$, camera intrinsics, and camera poses are stored in a training buffer. However, this strategy may sample many meaningless regions. To address this, FocusTune [9] projects the 3D point cloud of the scene onto the training images and performs uniform sampling around the projected points. While this approach better targets informative areas, it cannot guarantee consistent sampling across multiple images.
In contrast, we utilize the keypoint detection results from Section IV-B to select the top 1,000 patches with the highest confidence scores in each image, aligning with the number used by ACE. This approach ensures consistent patch selection across image sequences. It also ensures the same patches can be used during mapping and relocalization. We show a comparison of patch selection in Fig. 3.
Self-Loss
We follow the previous works [8, 10] to compute a reprojection loss to supervise the training of $f_{H}$:
$\mathcal{L}_{S,i}=\left\{\begin{array}[]{lcl}\tau(t)\tanh\left(\frac{e_{\pi}(\mathbf{p}_{i},\mathbf{P}_{i}^{w},\mathbf{T}_{wc})}{\tau(t)}\right)&\mbox{if}\;\mathbf{P}_{i}^{w}\in\mathcal{V}\\ \lVert\mathbf{P}_{i}^{w}-\overline{\mathbf{P}}_{i}^{w}\lVert&\mbox{otherwise},\end{array}\right.$ (6)
where $\mathbf{P}_{i}^{w}$ is the 3D scene point predicted by $f_{H}$ using $\mathbf{f}_{i}$ and $e_{\pi}(\mathbf{p}_{i},\mathbf{P}_{i}^{w},\mathbf{T}_{wc})$ is the reprojection error. $\mathcal{V}$ is a set of valid predictions that are between 0.1m and 1000m in front of the image plane and have a reprojection error of less than 1000px. $\scriptstyle{\overline{\mathbf{P}}_{i}^{w}}$ is the pseudo ground truth 3D point computed by the inverse projection of $\mathbf{p}_{i}$ with the camera pose $\mathbf{T}_{wc}$ and a fixed target depth at 10m. $\tau(t)$ is a dynamically adjusted scale between $\tau_{\min}=1$ and $\tau_{\max}=50$:
$\tau(t)=\tau_{\max}\sqrt{1-t^{2}}+\tau_{\min},$ (7)
where $t\in(0,1)$ represents the relative training progress.
For each image patch, $\mathcal{L}_{S,i}$ uses information from a single image to supervise the prediction of $\mathbf{P}_{i}^{w}$ for each $\mathbf{f}_{i}$, so we call it self-loss. However, in classical geometry, a single viewpoint constrains a scene point to lie along a line of sight rather than determining its precise 3D position. For instance, $\mathcal{L}_{S,i}$ would be zero if the predicted $\mathbf{P}_{i}^{w}$ falls anywhere on the inverse-projection line of $\mathbf{p}_{i}$. Therefore, SCR systems heavily rely on effective implicit triangulation, i.e., the neural network must consistently identify the same region across multiple images so that the scene point can be constrained by observations from different viewpoints. Achieving this is particularly challenging in complex or repetitive texture environments, as shown in [10].
Cross-Loss
To tackle the above challenge, we propose explicitly integrating multi-view constraints into the mapping process, rather than relying solely on the neural network to learn them implicitly. The core idea is to associate the selected patches across different images using feature-matching techniques [29]. This approach enables us to project a predicted 3D point onto multiple frames and leverage the resulting reprojection errors for more effective supervision.
Inspired by visual SLAM techniques, we select keyframes to ensure that matching patch pairs have varying baselines. A frame is chosen as a keyframe if the matched features with the previous keyframe are less than half of the detected features. For each matching pair $(\mathbf{p}_{k,j},\mathbf{p}_{i})$, where $\mathbf{p}_{k,j}$ is the matched point on the keyframe, we define a loss as:
$\mathcal{L}_{C,i}=\left\{\begin{array}[]{lcl}\tau(t)\tanh\left(\frac{e_{\pi}(\mathbf{p}_{k,j},\mathbf{P}_{i}^{w},\mathbf{T}_{wk})}{\tau(t)}\right)&\mbox{if}\;\mathbf{P}_{i}^{w}\in\mathcal{V}\\ \lVert\mathbf{P}_{i}^{w}-\overline{\mathbf{P}}_{k,j}^{w}\lVert&\mbox{otherwise},\end{array}\right.$ (8)
where $\mathbf{T}_{wk}$ is the pose of the keyframe and $\overline{\mathbf{P}}_{k,j}^{w}$ is the pseudo ground truth 3D point computed by the inverse projection of $\mathbf{p}_{k,j}$. Since $\mathcal{L}_{C,i}$ uses different frames to supervise the prediction of scene points, we call it cross-loss.
Training Details
For each image patch, we store the corresponding feature vector $\mathbf{f}_{i}$, the pixel $\mathbf{p}_{i}$, the camera intrinsic $\mathbf{K}$, the camera pose $\mathbf{T}_{wc}$, the pixel matched on the keyframe $\mathbf{p}_{k,j}$, and the keyframe pose $\mathbf{T}_{wk}$, in the training buffer. We shuffle these patches and select $\mathcal{N}_{b}=5120$ patches in each iteration to compute the loss:
$\mathcal{L}=\frac{1}{\mathcal{N}_{b}}\sum\left(\mathcal{L}_{S,i}+\lambda\mathcal{L}_{C,i}\right),$ (9)
where $\lambda=0.5$ if $(\mathbf{p}_{k,j},\mathbf{p}_{i})$ is an inlier match, otherwise $\lambda=0$. We train $f_{H}$ over 16 complete iterations using the AdamW optimizer [30] with a learning rate between $-45e\text{-}$ 4 and $-35e\text{-}$ 3 and a 1-cycle schedule [31].
TABLE I: Comparison on the 7-Scenes dataset. The relocalization recall is the proportion of test frames below a (1cm, 1) pose error.
Mapping Map FPS* Relocalization Recall (%)
[…]
    Running time of relocalization measured in frame per second (FPS).
IV-D Dual-Mode Relocalization
In this section, we introduce how to reuse the map for relocalization. Our system supports two relocalization modes. The first is single-mode, which estimates the camera pose using a single image. This mode is efficient and straightforward. As shown in Fig. 2, the framework generates a set of 2D keypoints and their corresponding predicted 3D scene points. Using these 2D-3D correspondences, we apply PnP and RANSAC algorithms to compute the optimal pose by maximizing the number of inliers. Like most existing SCR methods, our single-mode does not leverage the constraints between adjacent frames in an image sequence. However, in real-world robotic applications, the previous frame provides valuable prior information for pose estimation. Therefore, we also design a sequence-mode that utilizes sequential information to enhance the relocalization. The details of this mode are introduced in the following section.
Figure 4: In our sequence-based relocalization, the predicted 3D scene points are used to constrain the pose estimation of subsequent frames and updated with new observations.
Pose Prediction
In traditional visual odometry methods [32], local constraints are established by first triangulating 3D points from historical frames and then using these points to constrain the current pose. Typically, triangulation involves solving an $\mathbf{A}\mathbf{x}=\mathbf{b}$ problem through matrix decomposition, which can be computationally intensive. To simplify this process, we propose a more efficient approach. Our key insight is that if the reprojection errors of a predicted scene point across multiple historical frames remain small, the prediction can be considered accurate and used to constrain the current pose. Based on this idea, we maintain a set of reliably predicted scene points for pose estimation.
We denote two adjacent frames as $\mathcal{I}_{m}$ and $\mathcal{I}_{n}$, where $\mathcal{I}_{n}$ is the current frame and $\mathcal{I}_{m}$ is the previous frame. Let $\mathcal{Y}_{m}$ and $\mathcal{X}_{m}$ represent the maintained set of 3D scene points and their corresponding 2D points on $\mathcal{I}_{m}$, respectively. Using optical flow [33], we track $\mathcal{X}_{m}$ to obtain $\mathcal{\bar{X}}_{n}$ on $\mathcal{I}_{n}$, along with their corresponding scene points $\mathcal{\bar{Y}}_{n}$ (denoted with a bar to indicate pre-update values). The pose of $\mathcal{I}_{n}$ is then computed as:
$\mathbf{\bar{T}}_{wn}=\mathfrak{g}\left(\mathcal{\bar{X}}_{n},\mathcal{\bar{Y}}_{n}\right),$ (10)
where $\mathfrak{g}$ represent the PnP solver with RANSAC.
Pose Update
The pose estimation in Eq. 10 only uses 2D-3D correspondences of historical frames and feature tracking results, so $\mathbf{\bar{T}}_{wn}$ can be refined with new predictions from the network for $\mathcal{I}_{n}$. Let $\mathcal{\hat{X}}_{n}$ and $\mathcal{\hat{Z}}_{n}$ denote the newly detected keypoints and the corresponding 3D scene points predicted by the network. Since the set of 2D-3D correspondences $\{\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\}$ typically contains many outliers, we first use PnP with RANSAC to compute a pose $\mathbf{\hat{T}}_{wn}=\mathfrak{g}\left(\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\right)$, then fuse it with $\mathbf{\bar{T}}_{wn}$ at the pose level, rather than directly using a Kalman filter. We perform a simple weighted average on the manifold to obtain the updated pose:
$\mathbf{T}_{wn}=\exp\left(\left(\bar{w}_{n}\log(\mathbf{\bar{T}}_{wn})^{\vee}+\hat{w}_{n}\log(\mathbf{\hat{T}}_{wn})^{\vee}\right)^{\wedge}\right),$ (11)
where $\exp$ and $\log$ represent the exponential and logarithmic maps, respectively. The weights $\bar{w}_{n}$ and $\hat{w}_{n}$ are computed as:
$\bar{w}_{n}=\frac{\mathcal{\bar{N}}_{n}}{\mathcal{\bar{N}}_{n}+\mathcal{\hat{N}}_{n}},\quad\hat{w}_{n}=\frac{\mathcal{\hat{N}}_{n}}{\mathcal{\bar{N}}_{n}+\mathcal{\hat{N}}_{n}}$ (12)
where $\mathcal{\bar{N}}_{n}$ and $\mathcal{\hat{N}}_{n}$ are the number of inliers when estimating $\mathbf{\bar{T}}_{wn}$ and $\mathbf{\hat{T}}_{wn}$ using $\mathfrak{g}$.
Scene Point Update
The maintained scene point set also requires updating based on new observations from $\mathcal{I}_{n}$. Besides adding newly detected points and their corresponding predicted 3D points to $\{\mathcal{\bar{X}}_{n},\mathcal{\bar{Y}}_{n}\}$, we also update tracked scene points. For a scene point $\mathbf{\bar{P}}^{w}_{i}\in\mathcal{\bar{Y}}_{n}$ and its corresponding tracked 2D point $\mathbf{\bar{p}}_{i}^{n}\in\mathcal{\bar{X}}_{n}$, if the pair $(\mathbf{\bar{p}}_{i}^{n},\mathbf{\bar{P}}^{w}_{i})$ is an outlier in Eq. 10, it will be removed from $\mathcal{\bar{Y}}_{n}$ and $\mathcal{\bar{X}}_{n}$. Otherwise, $\mathbf{\bar{P}}^{w}_{i}$ is updated in a direction that reduces the reprojection error on $\mathcal{I}_{n}$. The update direction is illustrated in Fig. 4, where $\scriptstyle{P}$ and $\scriptstyle{p}$ represent $\mathbf{\bar{P}}^{w}_{i}$ and $\mathbf{\bar{p}}_{i}^{n}$ for simplicity. The inverse projection line of $\mathbf{\bar{p}}_{i}^{n}$ is denoted as $\scriptstyle{\overrightarrow{ND}}$, with $\scriptstyle{D}$ as the foot of the perpendicular. The coordinates of $\scriptstyle{D}$ are:
$\mathbf{\bar{D}}^{w}_{i}$ $=\left(\left(\mathbf{\bar{P}}^{w}_{i}-\mathbf{t}_{wn}\right)^{\top}\mathbf{v}_{nd}\right)\mathbf{v}_{nd}-\left(\mathbf{\bar{P}}^{w}_{i}-\mathbf{t}_{wn}\right),$ (13a)
$\mathbf{v}_{nd}$ $=\mathbf{R}_{wn}\frac{\mathbf{K}_{n}^{-1}\mathbf{\bar{p}}_{i}^{n}}{\|\mathbf{K}_{n}^{-1}\mathbf{\bar{p}}_{i}^{n}\|},$ (13b)
where $\mathbf{R}_{wn}$ and $\mathbf{t}_{wn}$ are the rotation and translation components of $\mathbf{T}_{wn}$, respectively. $\mathbf{v}_{nd}$ is the direction vector of $\scriptstyle{\overrightarrow{ND}}$. Then $\mathbf{\bar{P}}^{w}_{i}$ can be updated through:
$\mathbf{P}^{w}_{i}=\mathbf{\bar{P}}^{w}_{i}+\beta_{n}\left(\mathbf{\bar{D}}^{w}_{i}-\mathbf{\bar{P}}^{w}_{i}\right),$ (14)
where $\beta_{n}$ is a scale factor. As $\mathbf{P}^{w}_{i}$ is observed across more frames, the influence of each new observation should diminish, allowing $\mathbf{P}^{w}_{i}$ to stabilize gradually. Hence, we set $\beta_{n}=1/N_{P}$, where $N_{P}$ denotes the number of observations.
Figure 5: The recall rates in repetitive texture environments as the threshold varies from 0 to 1cm (left) and the scene images (right).
V Experimental Results
In this section, we present the experimental results. The experiments are performed on three standard datasets for SCR tasks: 7-Scenes [34], 12-Scenes [35], and Cambridge Landmarks [15]. These datasets provide training data for mapping and test data for relocalization. We evaluate the two relocalization modes introduced in Section IV-D. In the following sections, “Ours-Single” refers to the single-mode relocalization, while “Ours-Sequence” or “Ours” refers to the sequence-based relocalization. All experiments are conducted on an NVIDIA GeForce RTX 3090 GPU.
V-A Comparison with State-of-the-Art
7-Scenes
The 7-Scenes dataset comprises RGB-D images captured in seven indoor scenes using a Kinect RGB-D camera with a resolution of $640\times 480$. We train $f_{H}$ using the RGB frames and their ground-truth poses. Our approach is evaluated against several SOTA SCR systems and FM-based methods. Key metrics, including mapping time, map size, relocalization speed, and recall rate, are evaluated. The recall rate is defined as the percentage of test frames with pose errors below (1cm, 1). Results are summarized in Table I. Our single-mode relocalization achieves remarkable efficiency, running at 90, which is 1.6$\times$ faster than ACE, while also improving its recall by 6.4%. The sequence mode further delivers SOTA accuracy with high efficiency. By comparison, FocusTune and GLACE, both variants of ACE, have notable drawbacks: FocusTune relies on a 3D model and achieves only a modest accuracy gain (+3.9%), while GLACE sacrifices mapping and relocalization efficiency. In contrast, our system achieves a substantial recall improvement (+11%), while preserving ACE’s fast mapping time, compact map size, and relocalization speed.
The proposed method significantly enhances system performance in environments with repetitive textures. The “Stairs” sequence is captured on a staircase with numerous repetitive patterns, which poses a significant challenge for accurate relocalization. In this scenario, our single-mode and sequence-mode approaches improve the recall rate of ACE by 7.2% and 13.6%, respectively. This highlights the effectiveness of the sequence-based mapping and relocalization methods, contributing recall rate improvements of 8.9% and 4.8%, respectively. Fig. 5 illustrates scene images and the cumulative error distribution from 0 to 1cm for different systems under this challenging scenario.
TABLE II: Comparison on the 12-Scenes dataset. We report the percentage of test frames below a (1cm, 1) pose error.
Sequence DSAC* ACE FocusTune GLACE Ours
Apt1_kitchen 87.1 84.9 88.0 93.3 97.5
Apt1_living 91.3 89.5 85.2 92.9 95.9
Apt2_bed 89.3 83.6 88.5 88.5 90.2
Apt2_kitchen 83.0 84.3 90.0 93.0 90.9
Apt2_living 83.6 89.1 89.4 91.1 91.6
Apt2_luke 66.5 76.4 77.1 82.7 84.8
Office1_gates362 73.1 73.6 76.2 94.6 81.1
Office1_gates381 62.3 68.2 70.0 85.2 78.8
Office1_lounge 75.8 70.6 69.1 82.3 81.7
Office1_manolis 70.9 72.9 71.4 84.4 77.2
Office2_5a 62.6 60.2 54.1 68.6 73.8
Office2_5b 58.3 71.4 61.2 84.0 85.9
[…]
FPS 48 53 56 33 60
TABLE III: Comparison on the Cambridge Landmarks dataset. Metrics include the median pose error and the recall rate.
Sequence Median Error (cm / ) Recall (%)
ACE Ours $\Downarrow$ $\Uparrow$ ACE Ours $\Downarrow$ $\Uparrow$
GreatCourt 43 / 0.2 38 / 0.2 -5 / 0 6.7 9.5 +2.8
KingsCollege 28 / 0.4 25 / 0.4 -3 / -0 11.7 12.8 +1.1
OldHospital 31 / 0.6 27 / 0.6 -4 / 0 10.4 9.9 -0.5
ShopFacade 5 / 0.3 5 / 0.3 -0 / -0 85.4 86.4 +1.0
StMarysChurch 18 / 0.6 17 / 0.6 -1 / 0 20.2 23.0 +2.8
[…]
12-Scenes
The 12-Scenes dataset, collected in 12 scenes using an iPad color camera, provides RGB sequences at a resolution of $1296\times 968$. Like previous methods, we resize the image to $642\times 480$ during relocalization. We evaluate our system alongside several SOTA SCR approaches, with results summarized in Table II. Compared to the baseline (ACE), our methods improve relocalization recall by 8.6% and boost relocalization speed by 7. Although GLACE achieves slightly higher accuracy on this dataset, it comes at a significant cost: requiring 2.5$\times$ the mapping time, consuming 3$\times$ the memory for map storage, and operating at only half the relocalization speed of our approach. In Fig. 6, we compare the trajectory errors and recall rates of different algorithms for relocalization in two scenes.
Cambridge Landmarks
This dataset comprises five scenes from the old town of Cambridge, captured using a Google LG Nexus 5 smartphone. The captured videos were downsampled into low-frame-rate image sequences, leading to large motion gaps between consecutive frames. These gaps make feature tracking via optical flow challenging, so we evaluate only our single-mode relocalization on this dataset. The results, summarized in Table III, compare our system with ACE to highlight the benefits of incorporating sequential information during mapping. Metrics include the median pose error and the recall, defined as the percentage of test frames below a (10cm, 10°) pose error. Our methods reduce pose error by 3cm and improve recall by 1.4%, demonstrating the effectiveness of leveraging sequential information.
Figure 6: The relocalization trajectories of different SCR systems in two scenarios (top: Apt1_kitchen, bottom: Chess). We also report the recall rates, defined as the percentage of test frames below a (1cm, 1°) pose error.
TABLE IV: Comparison with other sequential SCR methods. We report the median pose error (cm / °) and the running speed.
Sequence Cai et al. [36] Tang et al. [37] KFNet [23] Ours
Chess 2.0 / 0.8 2.0 / 0.7 1.8 / 0.7 0.5 / 0.1
Fire 2.0 / 1.0 2.0 / 0.8 2.3 / 0.9 0.7 / 0.3
Heads 4.0 / 2.7 1.0 / 0.8 1.4 / 0.8 0.4 / 0.2
Office 3.0 / 0.8 3.0 / 0.8 2.5 / 0.7 0.9 / 0.2
Pumpkin 4.0 / 1.1 4.0 / 1.1 3.7 / 1.0 0.9 / 0.2
Kitchen 4.0 / 1.1 3.0 / 1.1 3.8 / 1.2 0.7 / 0.2
Stairs 18.0 / 3.9 4.0 / 1.6 3.3 / 0.9 2.4 / 0.7
[…]
FPS - 15 20 59
V-B Comparison with Sequential Methods
We further compare our method with other SCR approaches that leverage sequential information, using the 7-Scenes dataset for evaluation. The median pose errors and running speeds of each system are summarized in Table IV. Our system outperforms all others, achieving the best accuracy across all seven scenes with 67% lower translation error and 67% lower rotation error compared to the second-best method. This improvement likely stems from the fact that the other methods encode all pixels, including non-informative or misleading regions, which can dilute the network’s encoding capacity and hinder mapping accuracy. Additionally, [36] introduces an extra module for dense feature extraction, while [37] and KFNet [23] rely on dense matching, which compromises efficiency. In contrast, our method employs a lightweight head for detecting sparse features and efficiently tracks them using optical flow, striking a better balance between accuracy and computational cost.
TABLE V: Ablation study on the proposed keypoint detection head, sequence-based mapping, and sequence-based relocalization.
7-Scenes 12-Scenes
Recall (%) $\Uparrow$ Recall (%) $\Uparrow$
w/o KDH, SM, SR 55.2 77.1
w/o SM, SR 60.3 +5.1 80.4 +3.3
w/o SR 61.6 +6.4 81.4 +4.3
Ours (Full) 66.2 +11.0 85.7 +8.6
V-C Ablation Study
We conduct an ablation study on the 7-Scenes and 12-Scenes datasets to evaluate the contributions of the proposed keypoint detection head (KDH), sequence-based mapping (SM), and sequence-based relocalization (SR). By sequentially removing these three modules, we calculate the recall rate, defined as the percentage of test frames below a (1cm, 1°) pose error. The results, presented in Table V, clearly demonstrate the effectiveness of the proposed components. Since our system builds upon ACE, the configuration “w/o KDH, SM, SR” corresponds to the original ACE. On the 7-Scenes dataset, KDH, SM, and SR improve the recall rate by 5.1%, 1.3%, and 4.6%, respectively, achieving a total improvement of 11.0%. Similarly, on the 12-Scenes dataset, these methods lead to an 8.6% increase in recall rate, further validating their impact.
VI Conclusion
In this work, we propose an efficient and accurate SCR system. Our approach introduces a unified scene-agnostic module for both scene encoding and salient keypoint detection, enabling the system to prioritize encoding informative regions. To strengthen implicit triangulation, we associate keypoints across multiple images to effectively leverage sequential information. We further design two relocalization modes: the single-mode, which performs one-shot relocalization with high efficiency, and the sequence-mode, which incorporates temporal information to boost relocalization performance. Experimental results demonstrate that our methods significantly improve system performance, achieving remarkable accuracy and efficiency. Nevertheless, like other SCR systems, we use a CNN as the backbone, which limits scene point predictions to rely solely on information from a single patch. In future work, we plan to explore other architectures like ViT [38] to enhance SCR systems.

### v2

@@
-An Efficient Scene Coordinate Encoding and Relocalization Method
+Enhancing Scene Coordinate Regression with Efficient Keypoint Detection and Sequential Information
 Abstract
@@
 However, current SCR methods often face challenges in handling repetitive textures and meaningless areas due to their reliance on implicit triangulation.
-In this paper, we propose an efficient scene coordinate encoding and relocalization method.
-Compared with the existing SCR methods, we design a unified architecture for both scene encoding and salient keypoint detection, enabling our system to focus on encoding informative regions, thereby significantly enhancing efficiency.
-Additionally, we introduce a mechanism that leverages sequential information during both map encoding and relocalization, which strengthens implicit triangulation, particularly in repetitive texture environments.
-Comprehensive experiments conducted across indoor and outdoor datasets demonstrate that the proposed system outperforms other state-of-the-art (SOTA) SCR methods.
+In this paper, we propose an efficient and accurate SCR system.
+Compared to existing SCR methods, we propose a unified architecture for both scene encoding and salient keypoint detection, allowing our system to prioritize the encoding of informative regions.
+This design significantly improves computational efficiency.
+Additionally, we introduce a mechanism that utilizes sequential information during both mapping and relocalization.
+The proposed method enhances the implicit triangulation, especially in environments with repetitive textures.
+Comprehensive experiments conducted across indoor and outdoor datasets demonstrate that the proposed system outperforms state-of-the-art (SOTA) SCR methods.
 Our single-frame relocalization mode improves the recall rate of our baseline by 6.4% and increases the running speed from 56 to 90.
@@
 Visual relocalization is crucial for many applications such as mobile robots and augmented reality, offering a practical balance between cost and accuracy [1].
-Its core task involves estimating the camera pose using an existing map, which can be broadly divided into two categories: explicit maps and implicit maps.
+The core task involves estimating the camera pose using a pre-built map.
+The map can be generally classified into two categories: explicit maps and implicit maps.
 Explicit maps consist of a 3D point cloud combined with visual descriptors of 2D features [2, 3], while implicit maps encode the scene using a neural network [1, 4].
@@
 Recently, ACE [8] proposes a generalized convolutional neural network (CNN) to extract feature maps for scene encoding.
-Building on this pre-trained backbone, it utilizes a compact 4MB-sized multilayer perceptron (MLP) head as the map, enabling new scene encoding in five minutes only without requiring a 3D model or depth information.
+Building on this pre-trained backbone, ACE utilizes a compact 4MB-sized multilayer perceptron (MLP) head as the map, enabling new scene encoding in five minutes only without requiring a 3D model or depth information.
 This approach achieves state-of-the-art accuracy and efficiency.
@@
 Our system delivers accurate and efficient relocalization, while maintaining fast mapping and a compact map size, achieving state-of-the-art performance.
-Despite their advantages, SCR methods face several notable limitations.
+Despite their advantages, SCR methods still face several notable limitations.
 First, the presence of many texture-less and non-informative regions in the image significantly increases the map encoding errors [9].
@@
 Finally, the 6-DOF pose of the query image is estimated using geometric techniques, such as Perspective-n-Point (PnP) and bundle adjustment [13, 14].
-However, FM-based methods require storing 2D keypoints, visual descriptors, and a 3D point cloud, resulting in relatively large map sizes.
-For large-scale scenes, managing these maps demands significant memory resources.
-Furthermore, during relocalization, multiple candidate frames are typically retrieved for feature matching, which imposes computational overhead and reduces efficiency.
 II-B Absolute Pose Regression
-Absolute pose regression (APR) approaches use a DNN to regress camera poses from RGB images directly.
+Unlike FM-based methods, absolute pose regression (APR) approaches use a DNN to regress camera poses from RGB images directly.
 They have also garnered notable attention recently due to their good efficiency and simple architecture.
-PoseNet [15] is the first ARP solution which uses GoogLeNet [16] as the backbone.
-The following works explore different architectural designs to enhance ARP systems [17, 18].
-Other works leverage sequential information to improve APR accuracy by incorporating temporal filtering [19] or visual odometry [20].
-Recent ARP works utilize novel synthesis techniques to generate high-quality training data [21] or integrate them for fine-tuning [22].
-However, ARP approaches still fall short in accuracy compared to SCR and FM-based methods.
+PoseNet [15] is the first APR system.
+The subsequent works explore different architectural designs [16] or leverage sequential information [17] to enhance APR systems.
+Recent APR works utilize novel synthesis techniques to generate high-quality training data [18] or integrate them for fine-tuning [19].
+However, APR approaches still fall short in accuracy compared to SCR and FM-based methods.
 II-C Scene Coordinate Regression
-Different from ARP approaches, SCR methods regress the corresponding 3D coordinates in the scene for 2D pixels in the query image.
-Most existing SCR methods require depth information or a 3D model of the scene.
+Different from FM-based methods and APR approaches, SCR methods regress the corresponding 3D coordinates in the scene for 2D pixels in the query image.
+Most existing SCR methods require depth information or a 3D model of the scene [20, 21].
 Dong et al.
-[7] utilize a pre-learned feature extractor and scene region classifier to encode a scene with few-shot images, achieving efficient mapping.
-However, the additional modules, e.g., feature detection and hierarchical classification, limit its localization speed to only $5$ even on a powerful GPU.
+[7] use a pre-trained feature extractor and scene classifier for efficient few-shot mapping, but extra modules limit localization speed to $5$ even on a high-end GPU.
 Zhou et al.
-[23] propose KFNet.
-They leverage the sequential information by integrating the traditional Kalman filter [24] into the SCR pipeline.
+[22] introduce KFNet, integrating a Kalman filter [23] into the SCR pipeline to leverage sequential information.
 Bui et al.
-[25] propose D2S, which builds on a pre-trained sparse feature extraction network.
-They establish a mapping between feature descriptors and 3D scene points.
-DSAC++ [26] is the first SCR method that does not rely on depth or 3D model information.
-However, it requires a long training time.
-DSAC* [1] simplifies the DSAC++ training procedure by unifying two separate initialization steps, reducing the training time from 6 days to 15 hours.
-ACE [8] replaces the backbone of DSAC* with a pre-trained CNN, requiring only a new header to be trained when encoding a new scene, which greatly improves the mapping efficiency.
-This reduces the mapping time to 4 minutes and the map size to 4MB.
-FocusTune [9] and GLACE [10] are variations of ACE.
-FocusTune projects a 3D model onto training images to avoid encoding meaningless image regions, but this introduces the dependency on the 3D model.
-GLACE incorporates global image descriptors into the ACE pipeline to enhance the distinguishability of similar regions.
+[24] propose D2S, which maps sparse feature descriptors to 3D scene points using a pre-trained network.
+DSAC++ [25] is the first SCR method independent of depth or 3D models but requires long training.
+DSAC* [1] streamlines DSAC++ by unifying initialization steps, cutting training time from 6 days to 15 hours.
+ACE [8] further improves efficiency by using a pre-trained CNN, reducing mapping time to 4 minutes and map size to 4MB.
+FocusTune [9], GLACE [10], and EGFS [26] are all variants of ACE, each introducing unique enhancements.
+FocusTune refines ACE by projecting a 3D model onto images to filter irrelevant regions, but it depends on the 3D model.
+EGFS refines region selection through an iterative process using a segmentation model and reprojection errors.
+Meanwhile, GLACE enhances ACE by incorporating global image descriptors for better region distinction.
+Our system also builds on ACE.
+However, unlike these methods, we utilize keypoint detection to filter out uninformative regions and leverage sequential information to reduce visual aliasing.
+II-D Keypoint Detection
+Keypoint detection is essential for visual localization.
+Traditional methods like FAST [27] and ORB [28] rely on handcrafted detectors and descriptors, offering efficiency but often lacking robustness.
+In contrast, learning-based approaches, such as SuperPoint [29] and R2D2 [30], leverage deep networks to enhance repeatability and precision.
+Recently, detector-free methods like D2-Net [31] jointly learn detection and description, improving performance in challenging scenarios.
+To optimize efficiency, keypoint detection can leverage a shared backbone network in multi-task learning, enabling integration with tasks like local-global feature extraction [3] or joint point-line detection [6].
+Inspired by [3, 6], we use a shared backbone for both SCR and keypoint detection, improving system efficiency.
 III Preliminaries
@@
 The objective is to estimate the pose $\mathbf{T}_{wc}$ of a given image $I$, where $\mathbf{T}_{wc}$ denotes the transformation from the camera coordinate system to the world coordinate system.
-Given a set of 2D-3D correspondences $\mathcal{C}={(\mathbf{p}_{i},\mathbf{P}_{i}^{w})}$, where $\mathbf{p}_{i}$ represents a pixel in the image $I$ and $\mathbf{P}_{i}^{w}$ is its corresponding 3D point in the world coordinate system, the pose can be efficiently computed using PnP algorithms or iterative optimization techniques.
+Given a set of 2D-3D correspondences $\mathcal{C}={(\mathbf{p}_{i},\mathbf{P}_{i}^{w})}$, where $\mathbf{p}_{i}$ represents a pixel in the image $I$ and $\mathbf{P}_{i}^{w}$ is the corresponding 3D point in the world coordinate system, the pose can be efficiently computed using PnP algorithms or iterative optimization techniques.
 To obtain these correspondences $\mathcal{C}$, SCR methods aim to learn a mapping from 2D pixels to their corresponding 3D points:
@@
 2.
-It comprises two main components: a scene-agnostic module and a scene-specific module.
+Our system consists of a scene-agnostic module and a scene-specific module.
 The scene-agnostic module generates feature vectors and saliency scores for image patches, while the scene-specific module, designed as an MLP head with several 1D convolutional layers, maps these feature vectors to 3D scene points.
@@
 However, employing two separate deep neural networks would significantly reduce online efficiency.
-Inspired by [3, 6], we think that these two tasks can be accomplished using a shared backbone.
+Inspired by [3, 6], we accomplish these two tasks with a shared backbone.
 To achieve this, we extend the ACE backbone by adding a lightweight keypoint detection head (KDH).
 The KDH consists of just three convolutional layers, introducing trivial computational overhead.
-Its input is the feature map from the backbone, and its output is a tensor $\mathcal{T}_{K}\in\mathbb{R}^{65\times\frac{H}{8}\times\frac{W}{8}}$, where $H$ and $W$ denote the height and width of the input image, respectively.
+The KDH takes the feature map from the backbone as input and outputs a tensor $\mathcal{T}_{K}\in\mathbb{R}^{65\times\frac{H}{8}\times\frac{W}{8}}$, where $H$ and $W$ denote the height and width of the input image, respectively.
 The 65 channels correspond to an $8\times 8$ grid region and a dustbin channel representing the absence of keypoints.
 This tensor is processed through a softmax layer and resized into a $H\times W$ heatmap, which scores the saliency of the corresponding image pixels.
-Since the KDH, like the backbone $f_{B}$ in Eq.
-2, is also scene-agnostic, we combine them into a unified module denoted as $f_{S}$, with its parameters represented by $\mathbf{w}_{S}$.
+Since the KDH, similar to the backbone $f_{B}$ in Eq.
+2, is scene-agnostic, we integrate them into a unified module, denoted as $f_{S}$, with parameters represented by $\mathbf{w}_{S}$.
 Then the second equation in Eq.
@@
 In the first step, we follow the approach outlined in [8] to train the backbone $f_{B}$.
-In the second step, the parameters of $f_{B}$ are fixed, and we utilize distillation techniques to train the KDH.
-Theoretically, any keypoint detector could serve as the teacher network.
-For our setup, we use the prediction of SuperPoint [27] as supervision signals.
+In the second step, the parameters of $f_{B}$ are fixed, and the distillation techniques are utilized to train the KDH.
+We use the prediction of SuperPoint [29] as supervision signals as it is simple, efficient, and performs excellently.
 The loss function is defined as:
@@
 We also experimented with a multi-task learning approach [3] to train $f_{S}$, but found that the two-step pipeline produced better results.
-This is probably because the difficulty of the two tasks differs significantly: specifically, training a keypoint detector is much simpler than training a universal scene encoder.
-This disparity makes it challenging to balance the two tasks effectively.
+This may be attributed to the substantial difference in difficulty between these two tasks.
+Specifically, training a keypoint detector converges much faster than training a universal scene encoder.
+This disparity in task difficulty may lead to suboptimal training within a multitask network [33].
 Figure 3: Sampling strategies comparison.
-ACE randomly samples image patches.
-FocusTune performs uniform sampling around the projected points of the 3D model.
-We use keypoint detection to select repeatable image patches.
+ACE [8] randomly samples image patches, while FocusTune [9] performs uniform sampling around the projected points of the 3D model.
+In contrast, we employ keypoint detection to identify repeatable image patches.
 IV-C Sequence-Based Mapping
@@
 To address this, FocusTune [9] projects the 3D point cloud of the scene onto the training images and performs uniform sampling around the projected points.
-While this approach better targets informative areas, it cannot guarantee consistent sampling across multiple images.
+While this approach better targets informative areas, FocusTune cannot guarantee consistent sampling across multiple images.
 In contrast, we utilize the keypoint detection results from Section IV-B to select the top 1,000 patches with the highest confidence scores in each image, aligning with the number used by ACE.
-This approach ensures consistent patch selection across image sequences.
-It also ensures the same patches can be used during mapping and relocalization.
+This approach guarantees consistent patch selection across image sequences, and also ensures the same patches are utilized during both mapping and relocalization.
 We show a comparison of patch selection in Fig.
@@
 where $t\in(0,1)$ represents the relative training progress.
-For each image patch, $\mathcal{L}_{S,i}$ uses information from a single image to supervise the prediction of $\mathbf{P}_{i}^{w}$ for each $\mathbf{f}_{i}$, so we call it self-loss.
-However, in classical geometry, a single viewpoint constrains a scene point to lie along a line of sight rather than determining its precise 3D position.
-For instance, $\mathcal{L}_{S,i}$ would be zero if the predicted $\mathbf{P}_{i}^{w}$ falls anywhere on the inverse-projection line of $\mathbf{p}_{i}$.
-Therefore, SCR systems heavily rely on effective implicit triangulation, i.e., the neural network must consistently identify the same region across multiple images so that the scene point can be constrained by observations from different viewpoints.
-Achieving this is particularly challenging in complex or repetitive texture environments, as shown in [10].
+For each image patch, $\mathcal{L}_{S,i}$ supervises the prediction of $\mathbf{P}_{i}^{w}$ for each $\mathbf{f}_{i}$ using only a single image, which is why we refer to it as self-loss.
+However, in classical multi-view geometry, triangulating a point successfully requires at least two observations [10, 34].
+Therefore, SCR systems heavily rely on implicit triangulation, implying that the neural network must consistently recognize the same region across multiple images to ensure that scene points are constrained by observations from different viewpoints.
+This is particularly challenging under complex environments.
 Cross-Loss
@@
 For each image patch, we store the corresponding feature vector $\mathbf{f}_{i}$, the pixel $\mathbf{p}_{i}$, the camera intrinsic $\mathbf{K}$, the camera pose $\mathbf{T}_{wc}$, the pixel matched on the keyframe $\mathbf{p}_{k,j}$, and the keyframe pose $\mathbf{T}_{wk}$, in the training buffer.
-We shuffle these patches and select $\mathcal{N}_{b}=5120$ patches in each iteration to compute the loss:
+We follow ACE [8] to shuffle these patches and select $\mathcal{N}_{b}=5120$ patches in each iteration to compute the loss:
 $\mathcal{L}=\frac{1}{\mathcal{N}_{b}}\sum\left(\mathcal{L}_{S,i}+\lambda\mathcal{L}_{C,i}\right),$ (9)
@@
     Running time of relocalization measured in frame per second (FPS).
-IV-D Dual-Mode Relocalization
-In this section, we introduce how to reuse the map for relocalization.
+IV-D Relocalization
+IV-D1 Single-Mode
 Our system supports two relocalization modes.
-The first is single-mode, which estimates the camera pose using a single image.
-This mode is efficient and straightforward.
+The single-mode estimates the camera pose from a single image, offering efficiency and simplicity.
 As shown in Fig.
-2, the framework generates a set of 2D keypoints and their corresponding predicted 3D scene points.
-Using these 2D-3D correspondences, we apply PnP and RANSAC algorithms to compute the optimal pose by maximizing the number of inliers.
-Like most existing SCR methods, our single-mode does not leverage the constraints between adjacent frames in an image sequence.
+2, it detects 2D keypoints, predicts corresponding 3D scene points, and applies PnP with RANSAC to optimize the pose by maximizing inliers.
+Note that this mode also uses the map built with our sequence-based mapping method.
+IV-D2 Sequence-Mode
+Like most existing SCR methods, our single-mode does not leverage the constraints between adjacent frames in an image sequence during relocalization.
 However, in real-world robotic applications, the previous frame provides valuable prior information for pose estimation.
 Therefore, we also design a sequence-mode that utilizes sequential information to enhance the relocalization.
-The details of this mode are introduced in the following section.
+Details of this mode are in the following section.
 Figure 4: In our sequence-based relocalization, the predicted 3D scene points are used to constrain the pose estimation of subsequent frames and updated with new observations.
@@
 $\mathbf{\bar{T}}_{wn}=\mathfrak{g}\left(\mathcal{\bar{X}}_{n},\mathcal{\bar{Y}}_{n}\right),$ (10)
-where $\mathfrak{g}$ represent the PnP solver with RANSAC.
+where $\mathfrak{g}$ represents the PnP solver with RANSAC.
 Pose Update
@@
 Let $\mathcal{\hat{X}}_{n}$ and $\mathcal{\hat{Z}}_{n}$ denote the newly detected keypoints and the corresponding 3D scene points predicted by the network.
-Since the set of 2D-3D correspondences $\{\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\}$ typically contains many outliers, we first use PnP with RANSAC to compute a pose $\mathbf{\hat{T}}_{wn}=\mathfrak{g}\left(\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\right)$, then fuse it with $\mathbf{\bar{T}}_{wn}$ at the pose level, rather than directly using a Kalman filter.
+Since the set of 2D-3D correspondences $\{\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\}$ typically contains many outliers, we first use PnP with RANSAC to compute a pose $\mathbf{\hat{T}}_{wn}=\mathfrak{g}\left(\mathcal{\hat{X}}_{n},\mathcal{\hat{Z}}_{n}\right)$, then fuse $\mathbf{\hat{T}}_{wn}$ with $\mathbf{\bar{T}}_{wn}$ at the pose level, rather than directly using a Kalman filter.
 We perform a simple weighted average on the manifold to obtain the updated pose:
@@
 Besides adding newly detected points and their corresponding predicted 3D points to $\{\mathcal{\bar{X}}_{n},\mathcal{\bar{Y}}_{n}\}$, we also update tracked scene points.
-For a scene point $\mathbf{\bar{P}}^{w}_{i}\in\mathcal{\bar{Y}}_{n}$ and its corresponding tracked 2D point $\mathbf{\bar{p}}_{i}^{n}\in\mathcal{\bar{X}}_{n}$, if the pair $(\mathbf{\bar{p}}_{i}^{n},\mathbf{\bar{P}}^{w}_{i})$ is an outlier in Eq.
+For a scene point $\mathbf{\bar{P}}^{w}_{i}\in\mathcal{\bar{Y}}_{n}$ and the corresponding tracked 2D point $\mathbf{\bar{p}}_{i}^{n}\in\mathcal{\bar{X}}_{n}$, if the pair $(\mathbf{\bar{p}}_{i}^{n},\mathbf{\bar{P}}^{w}_{i})$ is an outlier in Eq.
 10, it will be removed from $\mathcal{\bar{Y}}_{n}$ and $\mathcal{\bar{X}}_{n}$.
@@
 7-Scenes
-The 7-Scenes dataset comprises RGB-D images captured in seven indoor scenes using a Kinect RGB-D camera with a resolution of $640\times 480$.
+The 7-Scenes dataset captures $640\times 480$ resolution image sequences in seven indoor scenes.
 We train $f_{H}$ using the RGB frames and their ground-truth poses.
@@
 Results are summarized in Table I.
-Our single-mode relocalization achieves remarkable efficiency, running at 90, which is 1.6$\times$ faster than ACE, while also improving its recall by 6.4%.
+Our single-mode relocalization achieves remarkable efficiency, running at 90, which is 1.6$\times$ faster than ACE, while also improving the recall by 6.4%.
 The sequence mode further delivers SOTA accuracy with high efficiency.
@@
 TABLE II: Comparison on the 12-Scenes dataset. We report the percentage of test frames below a (1cm, 1) pose error.
-Sequence DSAC* ACE FocusTune GLACE Ours
-Apt1_kitchen 87.1 84.9 88.0 93.3 97.5
-Apt1_living 91.3 89.5 85.2 92.9 95.9
-Apt2_bed 89.3 83.6 88.5 88.5 90.2
-Apt2_kitchen 83.0 84.3 90.0 93.0 90.9
-Apt2_living 83.6 89.1 89.4 91.1 91.6
-Apt2_luke 66.5 76.4 77.1 82.7 84.8
-Office1_gates362 73.1 73.6 76.2 94.6 81.1
-Office1_gates381 62.3 68.2 70.0 85.2 78.8
-Office1_lounge 75.8 70.6 69.1 82.3 81.7
-Office1_manolis 70.9 72.9 71.4 84.4 77.2
-Office2_5a 62.6 60.2 54.1 68.6 73.8
-Office2_5b 58.3 71.4 61.2 84.0 85.9
+Sequence DSAC*[1] ACE[8] Focus- GLACE[10] Ours
+Tune[9] Single Seq.
+Apt1_kitchen 87.1 84.9 88.0 93.3 95.2 97.5
+Apt1_living 91.3 89.5 85.2 92.9 90.9 95.9
+Apt2_bed 89.3 83.6 88.5 88.5 88.9 90.2
+Apt2_kitchen 83.0 84.3 90.0 93.0 84.3 90.9
+Apt2_living 83.6 89.1 89.4 91.1 92.5 91.6
+Apt2_luke 66.5 76.4 77.1 82.7 82.1 84.8
+Office1_gates362 73.1 73.6 76.2 94.6 78.0 81.1
+Office1_gates381 62.3 68.2 70.0 85.2 74.0 78.8
+Office1_lounge 75.8 70.6 69.1 82.3 72.2 81.7
+Office1_manolis 70.9 72.9 71.4 84.4 72.7 77.2
+Office2_5a 62.6 60.2 54.1 68.6 68.6 73.8
+Office2_5b 58.3 71.4 61.2 84.0 77.5 85.9
 […]
-FPS 48 53 56 33 60
+FPS 48 53 56 33 87 60
 TABLE III: Comparison on the Cambridge Landmarks dataset. Metrics include the median pose error and the recall rate.
-Sequence Median Error (cm / ) Recall (%)
-ACE Ours $\Downarrow$ $\Uparrow$ ACE Ours $\Downarrow$ $\Uparrow$
-GreatCourt 43 / 0.2 38 / 0.2 -5 / 0 6.7 9.5 +2.8
-KingsCollege 28 / 0.4 25 / 0.4 -3 / -0 11.7 12.8 +1.1
-OldHospital 31 / 0.6 27 / 0.6 -4 / 0 10.4 9.9 -0.5
-ShopFacade 5 / 0.3 5 / 0.3 -0 / -0 85.4 86.4 +1.0
-StMarysChurch 18 / 0.6 17 / 0.6 -1 / 0 20.2 23.0 +2.8
+Sequence Median Error (cm / ) Ensemble ($\times$4)
+HLoc[3] GLACE[10] ACE[8] Ours ACE[8] Ours
+GreatCourt 16/0.1 19/0.1 43/0.2 38/0.2 28/0.1 27/0.1
+KingsCollege 12/0.2 19/0.3 28/0.4 25/0.4 18/0.3 17/0.3
+OldHospital 15/0.3 17/0.4 31/0.6 27/0.6 25/0.5 24/0.5
+ShopFacade 4/0.2 4/0.2 5/0.3 5/0.3 5/0.3 5/0.3
+StMarysChurch 7/0.2 9/0.3 18/0.6 17/0.6 9/0.3 9/0.3
 […]
+  • In this table, “Ours" represents “Ours-Single."
 12-Scenes
-The 12-Scenes dataset, collected in 12 scenes using an iPad color camera, provides RGB sequences at a resolution of $1296\times 968$.
-Like previous methods, we resize the image to $642\times 480$ during relocalization.
-We evaluate our system alongside several SOTA SCR approaches, with results summarized in Table II.
+The 12-Scenes dataset provides $1296\times 968$ resolution image sequences across 12 scenes.
+Following prior works, we resize images to $642\times 480$ for relocalization.
+Table II presents the evaluation results against SOTA SCR methods.
 Compared to the baseline (ACE), our methods improve relocalization recall by 8.6% and boost relocalization speed by 7.
@@
 Cambridge Landmarks
-This dataset comprises five scenes from the old town of Cambridge, captured using a Google LG Nexus 5 smartphone.
+This dataset comprises five scenes from the old town of Cambridge.
 The captured videos were downsampled into low-frame-rate image sequences, leading to large motion gaps between consecutive frames.
 These gaps make feature tracking via optical flow challenging, so we evaluate only our single-mode relocalization on this dataset.
-The results, summarized in Table III, compare our system with ACE to highlight the benefits of incorporating sequential information during mapping.
-Metrics include the median pose error and the recall, defined as the percentage of test frames below a (10cm, 10°) pose error.
-Our methods reduce pose error by 3cm and improve recall by 1.4%, demonstrating the effectiveness of leveraging sequential information.
+The results, summarized in Table III, show that Ours-Single reduces the average error of ACE by 3cm, demonstrating the effectiveness of leveraging sequential information in enhancing mapping accuracy.
+However, when using the ensemble method, the improvement of Ours-Single over ACE becomes minimal.
+This may be because the clustering process disrupts the original image sequence order, resulting in very limited sequence information available for Ours-Single.
 Figure 6: The relocalization trajectories of different SCR systems in two scenarios (top: Apt1_kitchen, bottom: Chess).
 We also report the recall rates, defined as the percentage of test frames below a (1cm, 1°) pose error.
-TABLE IV: Comparison with other sequential SCR methods. We report the median pose error (cm / °) and the running speed.
-Sequence Cai et al. [36] Tang et al. [37] KFNet [23] Ours
-Chess 2.0 / 0.8 2.0 / 0.7 1.8 / 0.7 0.5 / 0.1
-Fire 2.0 / 1.0 2.0 / 0.8 2.3 / 0.9 0.7 / 0.3
-Heads 4.0 / 2.7 1.0 / 0.8 1.4 / 0.8 0.4 / 0.2
-Office 3.0 / 0.8 3.0 / 0.8 2.5 / 0.7 0.9 / 0.2
-Pumpkin 4.0 / 1.1 4.0 / 1.1 3.7 / 1.0 0.9 / 0.2
-Kitchen 4.0 / 1.1 3.0 / 1.1 3.8 / 1.2 0.7 / 0.2
-Stairs 18.0 / 3.9 4.0 / 1.6 3.3 / 0.9 2.4 / 0.7
+TABLE IV: Generalization validation.
+Methods 7-Scenes 12-Scenes Cambridge
+(%, $\uparrow$) (%, $\uparrow$) (cm / , $\downarrow$)
+ACE 55.2 77.1 25 / 0.4
+ACE + Ours 66.2 85.7 22 / 0.4
+GLACE 61.6 86.7 14 / 0.3
+GLACE + Ours 69.7 89.8 13 / 0.3
+  • 1
+    We report the percentage of test frames below a (1cm, 1) pose error on 7-Scenes and 12-Scenes, as well as the median pose errors on Cambridge Landmarks.
+V-B Generalization Validation
+Although our system builds upon ACE [8], the proposed methods are not dependent on the specific structure.
+To validate this, we also apply our methods to another SCR system, GLACE [10].
+As shown in Table IV, our methods yield improvements on both ACE and GLACE, where "ACE + Ours" and "GLACE + Ours" denote the application of our methods to ACE and GLACE, respectively.
+The results indicate consistent performance gains across all three datasets.
+However, the improvement on Cambridge Landmarks is relatively smaller, likely due to its limited sequential information compared to the 7-Scenes and 12-Scenes datasets, as discussed in Section V-A.
+Overall, these experimental results highlight both the effectiveness and generalization capability of the proposed methods.
+TABLE V: Ablation study on the proposed keypoint detection head, sequence-based mapping, and sequence-based relocalization.
 […]
-FPS - 15 20 59
-V-B Comparison with Sequential Methods
-We further compare our method with other SCR approaches that leverage sequential information, using the 7-Scenes dataset for evaluation.
-The median pose errors and running speeds of each system are summarized in Table IV.
-Our system outperforms all others, achieving the best accuracy across all seven scenes with 67% lower translation error and 67% lower rotation error compared to the second-best method.
-This improvement likely stems from the fact that the other methods encode all pixels, including non-informative or misleading regions, which can dilute the network’s encoding capacity and hinder mapping accuracy.
-Additionally, [36] introduces an extra module for dense feature extraction, while [37] and KFNet [23] rely on dense matching, which compromises efficiency.
-In contrast, our method employs a lightweight head for detecting sparse features and efficiently tracks them using optical flow, striking a better balance between accuracy and computational cost.
-TABLE V: Ablation study on the proposed keypoint detection head, sequence-based mapping, and sequence-based relocalization.
-7-Scenes 12-Scenes
-Recall (%) $\Uparrow$ Recall (%) $\Uparrow$
-w/o KDH, SM, SR 55.2 77.1
-w/o SM, SR 60.3 +5.1 80.4 +3.3
-w/o SR 61.6 +6.4 81.4 +4.3
-Ours (Full) 66.2 +11.0 85.7 +8.6
+black Methods 7-Scenes 12-Scenes Recall (%) $\Uparrow$ Recall (%) $\Uparrow$ w/o KDH, SM, SR 55.2 77.1 w/o SM, SR 60.3 +5.1 80.4 +3.3 w/o SR 61.6 +6.4 81.4 +4.3 Ours (Full) 66.2 +11.0 85.7 +8.6
+  • KDH, SM, and SR represent keypoint detection head, sequence-based mapping, and sequence-based relocalization, respectively.
 V-C Ablation Study
-We conduct an ablation study on the 7-Scenes and 12-Scenes datasets to evaluate the contributions of the proposed keypoint detection head (KDH), sequence-based mapping (SM), and sequence-based relocalization (SR).
-By sequentially removing these three modules, we calculate the recall rate, defined as the percentage of test frames below a (1cm, 1°) pose error.
-The results, presented in Table V, clearly demonstrate the effectiveness of the proposed components.
-Since our system builds upon ACE, the configuration “w/o KDH, SM, SR” corresponds to the original ACE.
-On the 7-Scenes dataset, KDH, SM, and SR improve the recall rate by 5.1%, 1.3%, and 4.6%, respectively, achieving a total improvement of 11.0%.
-Similarly, on the 12-Scenes dataset, these methods lead to an 8.6% increase in recall rate, further validating their impact.
+We perform an ablation study on the 7-Scenes and 12-Scenes datasets to assess the contributions of the proposed keypoint detection head (KDH), sequence-based mapping (SM), and sequence-based relocalization (SR).
+By sequentially removing these modules, we compute the recall rate, defined as the percentage of test frames with a pose error below (1cm, 1°).
+The results in Table V demonstrate the effectiveness of the proposed components.
+Since our system builds on ACE, the “w/o KDH, SM, SR” setting corresponds to ACE.
+On the 7-Scenes dataset, KDH, SM, and SR improve recall by 5.1%, 1.3%, and 4.6%, respectively, achieving a total improvement of 11.0%.
+On the 12-Scenes dataset, these methods yield an 8.6% gain, further validating their impact.
 VI Conclusion
-In this work, we propose an efficient and accurate SCR system.
-Our approach introduces a unified scene-agnostic module for both scene encoding and salient keypoint detection, enabling the system to prioritize encoding informative regions.
-To strengthen implicit triangulation, we associate keypoints across multiple images to effectively leverage sequential information.
-We further design two relocalization modes: the single-mode, which performs one-shot relocalization with high efficiency, and the sequence-mode, which incorporates temporal information to boost relocalization performance.
+In this work, we propose an efficient and accurate SCR system with a unified scene-agnostic module for scene encoding and keypoint detection, prioritizing informative regions.
+To enhance implicit triangulation, we associate keypoints across images to leverage sequential information.
+Additionally, we introduce two relocalization modes: a high-efficiency single-mode and a sequence-mode that integrates temporal cues for improved accuracy.
 Experimental results demonstrate that our methods significantly improve system performance, achieving remarkable accuracy and efficiency.
-Nevertheless, like other SCR systems, we use a CNN as the backbone, which limits scene point predictions to rely solely on information from a single patch.
-In future work, we plan to explore other architectures like ViT [38] to enhance SCR systems.
+Nevertheless, like other SCR systems, we use a CNN as the backbone, which limits scene point predictions to rely solely on information from a single local patch.
+In contrast, ViT [42] can capture long-range dependencies between image patches and leverage image global information to some extent.
+Therefore, for future work, we suggest training a scene-agnostic ViT encoder to extract feature vectors, enhancing the SCR system’s ability to leverage global image information.